help me solve 2/3-5/6=1/x
Answer by sdmmadam@yahoo.com(530)

solve 2/3-5/6=1/x ----(1)
Multiplying by 6x (which is the lcm of 2,3 and x)
2X(2x)-5X(x)= 1X(6)
4x-5x=6
-x = 6
That is x =-6
Answer: x = -6
Verification:The lcm on the left gives the left side=-1/6 =1/(-6)=1/x = RHS
Note: You may combine the two fractions on the LHS by finding the lcm and proceed too
[2X(2)-5X(1)]/6 = 1/x
(4-5)/6= 1/x
-1/6 =1/x
Recipracating (that is topsy turvying)
-6 = x
x=-6
Note: Reciprocating (p/q) means dividing 1 by (p/q) which becomes (q/p)